# Government Transparency Affects Innovation: Evidence from Wireless Products

### v1

@@ Abstract
Does government transparency affect innovation? I evaluate the launch of a government database with detailed technical information on the universe of wireless-enabled products on the U.S. market (N 347 thousand). The results show the launch approximately doubled the use of new technologies in the following ten years, an indicator of follow-on innovation. The increase affected both products in the same and new product classes, suggesting novelty; waned over several years, potentially due to an increase in secrecy and patenting; and boosted foreign more than U.S. domestic competitors. These results highlight the importance of information for private sector innovation.
JEL
O30, O33, O38
@@ 1 Introduction
Governments have long shaped the information environment in which firms innovate. Venetian medieval laws constrained the movement of glass workers to maintain a technological lead (Amato, 1997). 18th-century British parliamentary rewards system actively encouraged innovation disclosure, paying the inventor Thomas Lombe for exhibiting silk working machines in the Tower of London (Burrell and Kelly, 2015). The 1999 U.S. American Inventor’s Protection Act accelerated the disclosure of patent applications, and likely increased technology diffusion (Hegde et al., 2023). Today, support for open access innovation, calls for algorithmic transparency, as well as recent closures of U.S. government websites (Gotfredsen, 2025) likewise likely impact innovation.
The effects of transparency regulations on innovation may be large, given how heavily firms rely on secrecy in innovation. Surveys consistently find that firms rely on secrecy and lead time more than patenting in almost all industries across countries (Levin et al., 1987; Arundel, 2001; Cohen et al., 2000; Hall et al., 2014). In a 1994 U.S. survey—closest to the focus of this paper—Cohen et al. (2000, p. 33) finds that R&D laboratories in the communication equipment industry report patents effective as an appropriation mechanism for 26 percent of their products (midpoint mean, N 34), compared to 47 and 66 percent for secrecy and lead time. The role of patents is somewhat stronger in computers and electronic equipment industries, but all industries relevant to this analysis report secrecy as more effective than patents in product innovation. Any regulatory requirements to disclose information about new products may thus affect a major innovation appropriation mechanism.
@@
I study the effects of a 1998 government transparency shock that made this documentation available on the internet, and its effect on follow-on innovation by competitors and on the originator firm. I primarily rely on a difference-in-differences design for causal identification, which differentiates between follow-on innovation by competitors (who benefit from transparency) and by the originator firm (which does not). I supplement this approach with regression-discontinuity design for robustness and extensions to originator firm behavior.
@@ 2 Prior literature
Information is central to the economics of innovation, to a point where seminal work conceptualizes innovation as information (Nelson, 1959; Arrow, 1962; Stiglitz, 1999), and formal models focus on firms observing inventions of their competitors (Williams, 2017; Hall et al., 2014).
@@
Empirically, undisclosed innovation is difficult to observe as the proliferation of secrecy obfuscates the role of information in innovation. Most prior work relies on innovation observable in patents, in which firms voluntarily disclose their inventions in exchange for legal protection (Hall and Harhoff, 2012). Various study designs have leveraged variation in the costs of information access and the timing of disclosure to estimate the effects of patent disclosure on innovation. Cox (2019) document an increase in patenting after 1734 U.K. patent disclosure requirement (restricted to London, where patent descriptions were stored), and Furman et al. (2021) find a similar increase in U.S. locations where a patent library had opened between 1995 and 1997. Gross (2023) find a significant decrease in follow-on innovation due to delayed grant and publication of patents due to a Second World War patent secrecy program, and (Rassenfosse et al., 2024) find a similar decrease in patents subject to an analogous Cold War-era program. Multiple studies analyze the effects of the 1999 U.S. American Inventors Protection Act (AIPA), which accelerated patent application disclosure, and find positive effects on follow-on innovation and patent licensing (Hegde and Luo, 2018; Kim and Valentine, 2021; Baruffaldi and Simeth, 2020; Hegde et al., 2023).
@@
These findings provide consistent evidence of a positive effect of patent disclosure on follow-on innovation. However, they are limited to the minority of inventions that are (eventually) patented, which skews toward large firms in certain industries (Levin et al., 1987; Arundel, 2001; Cohen et al., 2000; Hall et al., 2014). Disclosure in patent documents is also a strategic choice and is complicated due to the willful infringement doctrine, which discourages inventors from reading patents (Sandrik, 2021). The effects of disclosure on innovation are thus confounded by market entry, self-selection, and strategic signaling.
@@
A fruitful approach to overcoming the limitations of patent data relies on firms having to disclose their products when entering a market. Moser (2005) relies on product data from 19th-century World’s Fairs, which arguably covered most high-tech inventions globally, (Argente et al., 2021) relies on barcode readers in stores to measure innovation in retail products. These studies confirm that most inventions were not patented, but did not evaluate the effects of disclosure on innovation. The open access movement provided some variation in science and innovation disclosure. Recent studies on open access mandates find a positive effect on follow-on innovation, relying mostly on patent and academic publications data (Bryan and Ozcan, 2021; Staudt, 2020), although its broader impact is disputed (Probst et al., 2023). Berkes and Nencka (2024) leverage variation in public library openings and find a sizable positive effect of library openings on future patenting. Closer to product market effects, Murray et al. (2016) find a positive effect of lower costs of access to genetically modified mice on entry of new researchers and exploration of more diverse research paths. Finally, in a simultaneous study, Kim et al. (2024) use a subset of the dataset used here to study the effects of open-source drivers on downstream Wifi router supply chains.
@@
This paper expands this literature by estimating the causal effects of government-mandated transparency on follow-on innovation outside of the patent system, inclusive of all products irrespective of intellectual property protections, and in a large high-tech industry. Broadly, it contributes to the literature on follow-on innovation and innovation spillovers, previously mostly limited to patent data (Bloom et al., 2013; Sampat and Williams, 2019), and to literature on the effects of government transparency, previously mostly limited to public-sector effects (Porumbescu et al., 2022; Corduneanu-Huci and Trlifaj, 2024).
@@ 3 Wireless products regulation
The empirical setup originates with the Federal Communication Commission (FCC), a U.S. government agency responsible for regulating wireless-enabled products since the 1934 adoption of Section 302 of the Communications Act, following an increase in incidents of interference of unrelated devices such as cash machines and TV receivers (Knapp and Wall, 1997).
@@
The documents submitted to the FCC in this regulatory process contain detailed information about the product and its producer. The purpose of this disclosure is for consumers, engineers, and regulatory agencies to be able to verify that marketed devices meet regulatory standards. Today, the documents are publicly accessible on the FCC’s website (this was not always the case, as I elaborate later).22endnote: 2Available at https://www.fcc.gov/oet/ea/fccid. They represent a detailed overview of the most recent developments in the universe of all wireless-capable products introduced to the U.S. market, which is representative of the frontier of wireless technological development globally. The website is a valuable resource for reporting on product introductions (Weatherbed and Lawler, 2024, e.g.), and the user manuals from this database are widely circulated on the internet. To my knowledge, the dataset is presented here for the first time in an academic study, apart from (Kim et al., 2024), who contemporaneously use it to analyze the effects of open-source drivers on Wifi router supply chains.
@@ 4 Dataset construction
I construct the dataset by web-scraping the universe of all product applications submitted to the FCC from 1981 to 2021 from FCC’s website. I start with firm identification based on a cleaned name and country and a list of all products and application dates of each firm.33endnote: 3I use Application ID instead of the better known FCC ID as a unit of analysis. About 20 percent of products contain more than one regulated wireless-capable device under different class or technology. To avoid a multi-level data structure, I perform all analysis on the Application ID level. For each product, I code its Class and the vector of frequencies it operates on. NewFrequency is a binary variable equal to 1 if the product uses a combination of frequencies not used by any prior product. The analysis primarily focuses on the subset of new products containing new frequencies.
@@
As the primary outcome variable, I count the future products that used the same combination of frequencies within 5 years after the introduction of the focal product, ForwardUse.44endnote: 4Robustness checks include three and seven year definitions. The count approximates the follow-on use of a narrowly defined technology after its introduction in the new product. The approach is akin to identifying innovative products based on new product characteristics (Argente and Yeh, 2022) or tracking future patent applications in a narrow patent class (Bell et al., 2018; Rigby, 2015; Hall et al., 2001). Compared to FCC product classes, which are broad and of which there are only about a hundred in the dataset, using frequency combinations generates thousands of unique narrow technology combinations in the period of interest. Based on a manual review of the data, products using the same set of frequencies are often closely related: for example, they may be two wireless crane controllers or two phones with identical wireless connectivity but different designs (see examples below).
@@
I count separately the future frequency use by the originator firm that applied for the first product (ForwardUse, Originator) and by other firms (ForwardUse, Treatment). There are about five times as many future uses by competitors as by the originator. Among competitors’ future use, I split the count between products by incumbents and new entrants (Incumbent, Entrant), products in the same class as the original product and in a different class (InClass, OutClass). Depending on whether the follow-on product was introduced by a country that joined the World Trade Organization (WTO) prior to 1998, I further split the count by domestic, foreign WTO and foreign non-WTO firms (Domestic, Foreign (Non-)WTO), and the combination of these splits.55endnote: 5Dates of accession taken from http://web.archive.org/web/20251021112700/https://www.wto.org/english/res_e/booksp_e/sli_e/4wtomembers.pdf.
@@
The FCC requires each applicant to submit documentation including external and internal photos, block diagram, schematics, parts list, technical description, list of operation frequencies, and a form listing product classification, applicant details, and product description. I code Post equal to 1 if these documents were made available online by the transparency shock. The applicant can claim confidentiality on some of the documents if they contain a trade secret or confidential information.66endnote: 6The FCC distinguishes between short- and long-term confidentiality. I disregard short-term confidentiality because it is highly correlated with long-term confidentiality and its effect is time-limited. Documents that may be designated as long-term confidential include: schematics, block diagrams, operational descriptions, parts lists, and tune-up information; and, in rare cases, internal photos and the user manual. The applicant must justify the request in a letter and pay a fee that, in 1998, amounted to 130 dollars. (Federal Communications Commission, 1998a). Applications classified for national security reasons are under a stricter regime, and are not available in this dataset here. I code Secrecy as a binary variable equal to 1 if the applicant claimed confidentiality. Even if the applicant claims confidentiality, some of the documents remain publicly available—an increase in information compared to the prior period. For this reason and to avoid overestimating the effects of transparency due to self-selection, even applications claiming secrecy are considered as being affected by the transparency shock in the post period.
@@
To control for confounding variables discussed below, I code Internet as a binary variable equal to 1 if the product uses frequency that is associated with 3G, 4G or wifi internet connectivity (details in Appendix LABEL:apx:sec:internet). I also code products that were subject to deregulation by mapping the deregulated FCC rules (details in Appendix LABEL:apx:sec:deregulated). I further match the dataset to USPTO / PatentsView data by firm cleaned name and country. I only consider U.S. patents even for foreign applicants, because the U.S. is the place of marketing of products under FCC regulation, and any potential infringement falls under the jurisdiction of the USPTO. RecentPatent is a binary variable equal to 1 if the firm applied for at least one patent within three years prior to the FCC application.
@@ 5 Transparency shock
In 1998, the FCC implemented several changes to streamline the authorization process. These changes were presented at a 1997 IEEE conference (Knapp and Wall, 1997), shown for public consultation (Federal Communications Commission, 1997a, b) and enacted in two orders (Federal Communications Commission, 1998a, b). As the main change of interest, the FCC digitalized the application process, and launched a publicly available website that details all products authorized by the FCC—the very website that I web-scraped in the data construction process.
@@
In addition to digitalizing the application process, the FCC made available all prior product authorizations going back to 1981. In a key difference, only basic technical information became available for products authorized prior to digitization, whereas products authorized since digitization include all exhibits in PDF format, except those claimed confidential by the applicant. The digitalization was voluntary in the first year of the announcement, but compliance built fast: just eight months after, virtually all applications were submitted digitally.77endnote: 7An alternative explanation of the fast uptake is that FCC staff was scanning remaining paper applications to PDFs, as they indicated in their decision (Federal Communications Commission, 1998b, p. 11,333). In either case, the effect was that from November 1998, exhibits of virtually all applications were available on the website.
@@
The policy shock did not change the public-access regime of the authorizations or the technical requirements. The information available online since the digitization had been available before through individual requests, as the FCC stated in response to concerns about confidentiality raised during the public consultation (Federal Communications Commission, 1998a). The option to claim secrecy on some documents had also been available since before the policy change, but few applicants used it. In 1993, five years prior to the changes, only nine percent of applications claimed secrecy. This increased dramatically in the years leading up to and after the policy shock. Five years after the change, in 2003, 71 percent of applicants claimed secrecy (see Figure LABEL:apx:fig:secrecy_patenting). Note that any reduction in the effects of the transparency shock due to this increase in secrecy would bias the results downward, as the empirical strategy treats all products in the post period as being exposed to the transparency shock, including those claiming confidentiality. Indeed, the results will show that as the use of secrecy increases following the transparency shock, the effects of the transparency shock diminish.
@@
The availability of data on secrecy and the increase in its use in the late 1990s initially sparked my interest in FCC transparency policies during this period. I conceptualize the transparency policy change as a lowering of the costs of information about products regulated by the FCC, and estimate the effects of this cost reduction on the follow-on use of technologies described in the submitted documents. Detailed information on the most recent wireless-capable products became instantly available for anyone with internet connectivity, right from the moment of marketing. The simultaneous increase in the use of secrecy, patenting, and other appropriation mechanisms (Hall et al., 2014), could signify a strategic response of firms that moderated the effect of the information availability over time. Next, I estimate the causal effect of the transparency shock on follow-on innovation, and provide some evidence on the strategic responses. Figure 3 presents the timeline of events, including several related changes during this period that I outline in the next section.
@@
5.1 Empirical context
Several other changes occurred during the relevant time period that contextualize the transparency shock, and risk conflating the estimates of its effects. The founding of the WTO in 1995 and the accession of various countries in the subsequent decade establishes a potential confounder. The trade liberalization associated with the WTO may have had various effects: it may have increased the introduction of foreign follow-on products by both the originator and competitor, increased originator product introduction by non-U.S. firms, or increased product introduction as a follow-on to a U.S. products. These changes are not time invariant, but the two empirical strategies should limit their impact and elucidate the robustness of the results.
In the difference-in-differences design, follow-on products introduced by both the originator (control) and the competitor (treatment) would be affected similarly by the time-variant changes. A potential issue arises if the originator and competitor are located in countries affected differently by trade liberalization. Sub-population results in table4 elucidate these cross-country patterns. For the most conservative estimation, it estimates follow-on innovation by competitors in the same country as the originator, including on a subset of U.S. originators only (column 3). It also shows results of U.S. competitors following-on a non-U.S. originator (column 4), where the impact of trade liberalization would bias the estimates downwards. The table further splits foreign competitor products by early (prior to 1998) WTO membership, and isolates follow-on products from China and Taiwan (columns 2 and 3). As most countries acceded to the WTO in 1995, any major change in follow-on use of new frequency combinations should be observable in the pre-period. The dynamic treatment plot in figure 4 shows few deviations in pre-trends since three years prior to the transparency shock, which corresponds to the founding of WTO.
The regression discontinuity design puts the most weight on observations near the cutt-off date. This lays two to three years after most countries joined the WTO in 1995 and 1996, and three to four years prior to the accession of China and Taiwan in 2001 and 2002. The second event is of particular concern, as it was associated with an increase in imports into the U.S. Table LABEL:apx:tab:diff_no_china_taiwan shows main results excluding firms from China and Taiwan from the analysis, both as originators and as competitors. All these results, discussed in detail later, show significant, if reduced, effect on follow-on product introduction by domestic and early WTO members.
Domestically, the Electronic Freedom of Information Act of 1996 mandated government agencies to publish specific records electronically by the end of 1999 (U.S.C., 1996a). I did not find any changes in FCC policies beyond the transparency changes detailed above. The Telecommunications Act of 1996 significantly revised the original 1934 law, focusing on enabling market entry in internet telecommunications services markets and increasing interoperability and standardization. It did not directly change the technical requirements in the product authorization process, but an amendment to Section 302 enabled the delegation of equipment testing and certification to private laboratories (U.S.C., 1996b; Emeritz, 1996). This meant that the mandatory testing documentation firms submit in their authorization applications became more often provided by an external testing provider. The causal identification strategies outlined below are robust to the effects of these changes, as products applied for in the treatment and control groups and in the pre- and post-periods would be affected equally by this change.
The American Inventor’s Protection Act (AIPA) of 1999 accelerated disclosure of patent applications. Prior literature found a positive effect on follow-on innovation which may have spilled-over to products that use patented technologies in the dataset (Hegde and Luo, 2018; Kim and Valentine, 2021; Baruffaldi and Simeth, 2020; Hegde et al., 2023). I control for this effect with a binary variable AIPA equal 1 for applications submitted by applicants with recent patents on or after its effective date of November 29, 2000.
A second set of changes the FCC implemented in 1998 changed the regulatory regime of several product categories, mainly by deregulating some products that do not actively emit electromagnetic signal (Federal Communications Commission, 1997a, 1998a). The deregulated products no longer had to acquire prior authorization from the FCC unless the firm submitted an application voluntarily. This implies that such products would stop appearing in the dataset. To avoid conflation from a selection bias stemming from the co-occurrence of the two events, I mapped regulatory changes to product categories and excluded all deregulated product from the analysis, including those submitted prior to the policy change. See Appendix LABEL:apx:sec:deregulated for details and validation of the process.
The period of interest also saw the entry of the first wireless internet-connected devices. To control for the effects of this development, I map the frequencies of 3G, 4G and Wifi technologies, and code Internet as a binary variable equal 1 if the product had internet capability. Appendix LABEL:apx:sec:internet lists these frequencies and details the process.

### v2

# Government Transparency and Innovation: Evidence from Wireless Products

@@ Abstract
Does government transparency affect innovation? I analyze the launch of a government database with detailed technical information on the universe of wireless-enabled products on the U.S. market (N 347 thousand). The results show that following the launch, the use of new technologies approximately doubled in the following ten years. The increase affected both products in the same and new product classes, suggesting novelty; waned over several years, potentially due to an increase in secrecy and patenting; and boosted foreign more than U.S. domestic competitors. These results highlight the importance of the information environment for private sector innovation.
Keywords:
innovation, transparency, competition, secrecy, government
@@ 1 Introduction
Governments have long shaped the information environment in which firms innovate. Venetian medieval laws constrained the movement of glass workers to maintain a technological lead (amato_island_1997). 18th-century British parliamentary rewards system actively encouraged innovation disclosure, paying the inventor Thomas Lombe for exhibiting silk working machines in the Tower of London (burrell_parliamentary_2015). The 1999 U.S. American Inventor’s Protection Act accelerated the disclosure of patent applications, and likely increased technology diffusion (hegde_patent_2023). Today, support for open access innovation, calls for algorithmic transparency, as well as recent closures of U.S. government websites (gotfredsen_fighting_2025) likewise likely impact innovation.
The effects of transparency regulations on innovation may be large, given how heavily firms rely on secrecy in innovation. Surveys consistently find that firms rely on secrecy and lead time more than patenting in almost all industries across countries (levin_appropriating_1987; arundel_relative_2001; cohen_protecting_2000; hall_choice_2014). In a 1994 U.S. survey—closest to the focus of this paper—cohen_protecting_2000 finds that R&D laboratories in the communication equipment industry report patents effective as an appropriation mechanism for 26 percent of their products (midpoint mean, N 34), compared to 47 and 66 percent for secrecy and lead time. The role of patents is somewhat stronger in computers and electronic equipment industries, but all industries relevant to this analysis report secrecy as more effective than patents in product innovation. Any regulatory requirements to disclose information about new products may thus affect a major innovation appropriation mechanism.
@@
I study the effects of a 1998 government transparency shock that made this documentation available on the internet, and its effect on follow-on innovation by competitors and on the originator firm. I primarily rely on a difference-in-differences design, which differentiates between follow-on innovation by competitors (who benefit from transparency) and by the originator firm (which does not). I supplement this approach with regression-discontinuity design for robustness and extensions to originator firm behavior.
@@ 2 Prior literature
Information is central to the economics of innovation, to a point where seminal work conceptualizes innovation as information (richard_r._nelson_simple_1959; arrow_economic_1962; kaul_knowledge_1999), and formal models focus on firms observing inventions of their competitors (williams_how_2017; hall_choice_2014).
@@
Empirically, undisclosed innovation is difficult to observe as the proliferation of secrecy obfuscates the role of information in innovation. Most prior work relies on innovation observable in patents, in which firms voluntarily disclose their inventions in exchange for legal protection (hall_recent_2012). Various study designs have leveraged variation in the costs of information access and the timing of disclosure to estimate the effects of patent disclosure on innovation. cox_patent_2019 document an increase in patenting after 1734 U.K. patent disclosure requirement (restricted to London, where patent descriptions were stored), and furman_disclosure_2021 find a similar increase in U.S. locations where a patent library had opened between 1995 and 1997. gross_hidden_2023 find a significant decrease in follow-on innovation due to delayed grant and publication of patents due to a Second World War patent secrecy program, and (rassenfosse_patents_2024) find a similar decrease in patents subject to an analogous Cold War-era program. Multiple studies analyze the effects of the 1999 U.S. American Inventors Protection Act (AIPA), which accelerated patent application disclosure, and find positive effects on follow-on innovation and patent licensing (hegde_patent_2018; kim_innovation_2021; baruffaldi_patents_2020; hegde_patent_2023).
@@
These findings provide consistent evidence of a positive effect of patent disclosure on follow-on innovation. However, they are limited to the minority of inventions that are (eventually) patented, which skews toward large firms in certain industries (levin_appropriating_1987; arundel_relative_2001; cohen_protecting_2000; hall_choice_2014). Disclosure in patent documents is also a strategic choice and is complicated due to the willful infringement doctrine, which discourages inventors from reading patents (sandrik_empirical_2021). The effects of disclosure on innovation are thus confounded by market entry, self-selection, and strategic signaling.
@@
A fruitful approach to overcoming the limitations of patent data relies on firms having to disclose their products when entering a market. moser_how_2005 relies on product data from 19th-century World’s Fairs, which arguably covered most high-tech inventions globally, (argente_patents_2021) relies on barcode readers in stores to measure innovation in retail products. These studies confirm that most inventions were not patented, but did not evaluate the effects of disclosure on innovation. The open access movement provided some variation in science and innovation disclosure. Recent studies on open access mandates find a positive effect on follow-on innovation, relying mostly on patent and academic publications data (bryan_impact_2021; staudt_mandating_2020), although its broader impact is disputed (probst_impact_2023). berkes_knowledge_2024 leverage variation in public library openings and find a sizable positive effect of library openings on future patenting. Closer to product market effects, murray_mice_2016 find a positive effect of lower costs of access to genetically modified mice on entry of new researchers and exploration of more diverse research paths. Finally, in a simultaneous study, kim_open_2024 use a subset of the dataset used here to study the effects of open-source drivers on downstream Wifi router supply chains.
@@
This paper expands this literature by analyzing the effects of government-mandated transparency on follow-on innovation outside of the patent system, inclusive of all products irrespective of intellectual property protections, and in a large high-tech industry. Broadly, it contributes to the literature on follow-on innovation and innovation spillovers, previously mostly limited to patent data (bloom_identifying_2013; sampat_how_2019), and to literature on the effects of government transparency, previously mostly limited to public-sector effects (porumbescu_government_2022; corduneanu-huci_freedom_2024).
@@ 3 Wireless products regulation
The empirical setup originates with the Federal Communication Commission (FCC), a U.S. government agency responsible for regulating wireless-enabled products since the 1934 adoption of Section 302 of the Communications Act, following an increase in incidents of interference of unrelated devices such as cash machines and TV receivers (knapp_considerations_1997).
@@
The documents submitted to the FCC in this regulatory process contain detailed information about the product and its producer. The purpose of this disclosure is for consumers, engineers, and regulatory agencies to be able to verify that marketed devices meet regulatory standards. Today, the documents are publicly accessible on the FCC’s website (this was not always the case, as I elaborate later). [^2] Available at https://www.fcc.gov/oet/ea/fccid. They represent a detailed overview of the most recent developments in the universe of all wireless-capable products introduced to the U.S. market, which is representative of the frontier of wireless technological development globally. The website is a valuable resource for reporting on product introductions (weatherbed_whats_2024, e.g.), and the user manuals from this database are widely circulated on the internet. To my knowledge, the dataset is presented here for the first time in an academic study, apart from (kim_open_2024), who contemporaneously use it to analyze the effects of open-source drivers on Wifi router supply chains.
@@ 4 Dataset construction
I construct the dataset by web-scraping the universe of all product applications submitted to the FCC from 1981 to 2021 from FCC’s website. I start with firm identification based on a cleaned name and country and a list of all products and application dates of each firm. [^3] I use Application ID instead of the better known FCC ID as a unit of analysis. About 20 percent of products contain more than one regulated wireless-capable device under different class or technology. To avoid a multi-level data structure, I perform all analysis on the Application ID level. For each product, I code its Class and the vector of frequencies it operates on. NewFrequency is a binary variable equal to 1 if the product uses a combination of frequencies not used by any prior product. The analysis primarily focuses on the subset of new products containing new frequencies.
@@
As the primary outcome variable, I count the future products that used the same combination of frequencies within 5 years after the introduction of the focal product, ForwardUse. [^4] Robustness checks include three and seven year definitions. The count approximates the follow-on use of a narrowly defined technology after its introduction in the new product. The approach is akin to identifying innovative products based on new product characteristics (argente_product_2022) or tracking future patent applications in a narrow patent class (bell_who_2018; rigby_technological_2015; hall_nber_2001). Compared to FCC product classes, which are broad and of which there are only about a hundred in the dataset, using frequency combinations generates thousands of unique narrow technology combinations in the period of interest. Based on a manual review of the data, products using the same set of frequencies are often closely related: for example, they may be two wireless crane controllers or two phones with identical wireless connectivity but different designs (see examples below).
@@
I count separately the future frequency use by the originator firm that applied for the first product (ForwardUse, Originator) and by other firms (ForwardUse, Treatment). There are about five times as many future uses by competitors as by the originator. Among competitors’ future use, I split the count between products by incumbents and new entrants (Incumbent, Entrant), products in the same class as the original product and in a different class (InClass, OutClass). Depending on whether the follow-on product was introduced by a country that joined the World Trade Organization (WTO) prior to 1998, I further split the count by domestic, foreign WTO and foreign non-WTO firms (Domestic, Foreign (Non-)WTO), and the combination of these splits. [^5] Dates of accession taken from http://web.archive.org/web/20251021112700/https://www.wto.org/english/res_e/booksp_e/sli_e/4wtomembers.pdf.
@@
The FCC requires each applicant to submit documentation including external and internal photos, block diagram, schematics, parts list, technical description, list of operation frequencies, and a form listing product classification, applicant details, and product description. I code Post equal to 1 if these documents were made available online by the transparency shock. The applicant can claim confidentiality on some of the documents if they contain a trade secret or confidential information. [^6] The FCC distinguishes between short- and long-term confidentiality. I disregard short-term confidentiality because it is highly correlated with long-term confidentiality and its effect is time-limited. Documents that may be designated as long-term confidential include: schematics, block diagrams, operational descriptions, parts lists, and tune-up information; and, in rare cases, internal photos and the user manual. The applicant must justify the request in a letter and pay a fee that, in 1998, amounted to 130 dollars. (federal_communications_commission_fcc_1998). Applications classified for national security reasons are under a stricter regime, and are not available in this dataset here. I code Secrecy as a binary variable equal to 1 if the applicant claimed confidentiality. Even if the applicant claims confidentiality, some of the documents remain publicly available—an increase in information compared to the prior period. For this reason and to avoid overestimating the effects of transparency due to self-selection, even applications claiming secrecy are considered as being affected by the transparency shock in the post period.
@@
To control for confounding variables discussed below, I code Internet as a binary variable equal to 1 if the product uses frequency that is associated with 3G, 4G or wifi internet connectivity (details in Supplemental Appendix LABEL:apx:sec:internet). I also code products that were subject to deregulation by mapping the deregulated FCC rules (details in Supplemental Appendix LABEL:apx:sec:deregulated). I further match the dataset to USPTO / PatentsView data by firm cleaned name and country. I only consider U.S. patents even for foreign applicants, because the U.S. is the place of marketing of products under FCC regulation, and any potential infringement falls under the jurisdiction of the USPTO. RecentPatent is a binary variable equal to 1 if the firm applied for at least one patent within three years prior to the FCC application.
@@ 5 Transparency shock
In 1998, the FCC implemented several changes to streamline the authorization process. These changes were presented at a 1997 IEEE conference (knapp_considerations_1997), shown for public consultation (federal_communications_commission_fcc_1997; federal_communications_commission_fcc_1997-1) and enacted in two orders (federal_communications_commission_fcc_1998; federal_communications_commission_fcc_1998-1). As the main change of interest, the FCC digitalized the application process, and launched a publicly available website that details all products authorized by the FCC—the very website that I web-scraped in the data construction process.
@@
In addition to digitalizing the application process, the FCC made available all prior product authorizations going back to 1981. In a key difference, only basic technical information became available for products authorized prior to digitization, whereas products authorized since digitization include all exhibits in PDF format, except those claimed confidential by the applicant. The digitalization was voluntary in the first year of the announcement, but compliance built fast: just eight months after, virtually all applications were submitted digitally. [^7] An alternative explanation of the fast uptake is that FCC staff was scanning remaining paper applications to PDFs, as they indicated in their decision (federal_communications_commission_fcc_1998-1, p. 11,333). In either case, the effect was that from November 1998, exhibits of virtually all applications were available on the website.
@@
The policy shock did not change the public-access regime of the authorizations or the technical requirements. The information available online since the digitization had been available before through individual requests, as the FCC stated in response to concerns about confidentiality raised during the public consultation (federal_communications_commission_fcc_1998). The option to claim secrecy on some documents had also been available since before the policy change, but few applicants used it. In 1993, five years prior to the changes, only nine percent of applications claimed secrecy. This increased dramatically in the years leading up to and after the policy shock. Five years after the change, in 2003, 71 percent of applicants claimed secrecy (see Figure LABEL:apx:fig:secrecy_patenting). Note that any reduction in the effects of the transparency shock due to this increase in secrecy would bias the results downward, as the empirical strategy treats all products in the post period as being exposed to the transparency shock, including those claiming confidentiality. Indeed, the results will show that as the use of secrecy increases following the transparency shock, the effects of the transparency shock diminish.
@@
The availability of data on secrecy and the increase in its use in the late 1990s initially sparked my interest in FCC transparency policies during this period. I conceptualize the transparency policy change as a lowering of the costs of information about products regulated by the FCC, and estimate the effects of this cost reduction on the follow-on use of technologies described in the submitted documents. Detailed information on the most recent wireless-capable products became instantly available for anyone with internet connectivity, right from the moment of marketing. The simultaneous increase in the use of secrecy, patenting, and other appropriation mechanisms (hall_choice_2014), could signify a strategic response of firms that moderated the effect of the information availability over time. Next, I estimate the increase in follow-on innovation following the transparency shock, and provide some evidence on the strategic responses. Figure 3 presents the timeline of events, including several related changes during this period that I outline in the next section.